OpenStudy (gg):

find limx^(1/(1-e^-x) when x --> 0

OpenStudy (gg):

\[\lim_{x \rightarrow 0}x ^{1/(1-e ^{-x})}\]

OpenStudy (gg):

\[\lim_{x \rightarrow 0}x ^{1/\ln (1-e ^{-x})}\]

OpenStudy (gg):

the last one is correct

OpenStudy (anonymous):

Take it piece by piece\[e ^{0}\]is what?

OpenStudy (gg):

it's 1

OpenStudy (anonymous):

\[\ln (1-1)\]=What is \[\ln 0\]It is a trick question.

OpenStudy (gg):

\[-\infty\]

OpenStudy (gg):

?

OpenStudy (anonymous):

Great. I never got that right each time my teacher asked. You are ahead of the game. So put it together, what is the answer?

OpenStudy (gg):

I got x^0, but x goes to 0, so I got 0^0

OpenStudy (anonymous):

Actually 1 over 0 to the power of infinity, I think. Either way our little game ended in a indeterminate form. This is one where you set y equal to your original thingy. Take ln y and take ln of your thingy. You ever did any of those?

OpenStudy (gg):

I don't understand what do you want to say, I am not so good with English. Can you just write solution?

OpenStudy (anonymous):

\[\ln y =\ln x ^{1/\ln(1-e ^{-x)}}\]

OpenStudy (gg):

and now I have to find limes of this? That's all? :)

OpenStudy (anonymous):

Take lim of each side

OpenStudy (gg):

ok :) thank you very much :)

OpenStudy (anonymous):

I forgot: L'hopital

OpenStudy (gg):

what L'hopital ? :)

OpenStudy (anonymous):

L'hopital = 1) derivative, 2) limit. Does not work here.